OpenStudy (anonymous):

Factor Completely. If the equation is prime, state this. 10r^2-29r+10

OpenStudy (anonymous):

(r-2.5)(r-0.4)

OpenStudy (anonymous):

How do you work it out though That is my main problem

OpenStudy (anonymous):

if you use the quadratic formula you can find out what r is. so for example if r=1 or 2 you know that (r-1)(r-2) will give you the answer right because: (r-1)(r-2) = 0 r-1 = 0 or r-2 = 0 r = 1 or r = 2 so you use the same logic for this question and you can use the quadratic formula to find out that r = 5/2 or r = 2/5

OpenStudy (anonymous):

do you understand..?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

sorry i'm really not sure how to explain then ><"